# From Tokens to Words: On the Inner Lexicon of LLMs

### v2

@@ Abstract
Natural language is composed of words, but modern LLMs process sub-words as input. A natural question raised by this discrepancy is whether LLMs encode words internally, and if so how. We present evidence that LLMs engage in an intrinsic detokenization process, where sub-word sequences are combined into coherent word representations. Our experiments show that this process takes place primarily within the early and middle layers of the model. They also show that it is robust to non-morphemic splits, typos and perhaps importantly—to out-of-vocabulary words: when feeding the inner representation of such words to the model as input vectors, it can “understand” them despite never seeing them during training. Our findings suggest that LLMs maintain a latent vocabulary beyond the tokenizer’s scope. These insights provide a practical, finetuning-free application for expanding the vocabulary of pre-trained models. By enabling the addition of new vocabulary words, we reduce input length and inference iterations, which reduces both space and model latency, with little to no loss in model accuracy.
@@ 1 Introduction
Large language models (LLMs) rely heavily on tokenization methods such as byte-pair encoding (BPE; Sennrich et al., 2016). Such methods often split words into multiple tokens, potentially disrupting their morphological structure (Batsuren et al., 2024). [^1] For instance, the word “unhappiness” might be tokenized into “un,” “h,” and “appiness” (see Fig. 1). Typos and other perturbations can also lead to large variations in the tokens that represent a word (Kaushal & Mahowald, 2022). Nonetheless, LLMs exhibit a remarkable ability to recover word meaning (Cao et al., 2023). This raises important questions about how models internally compose meaningful word representations from tokens, a process referred to as detokenization (Elhage et al., 2022; Gurnee et al., 2023).
@@
In this work, we seek to understand the detokenization mechanism in LLMs. We consider two cases: words that are not part of the model BPE vocabulary, and are thus split into multiple sub-words; and single-token, in-vocabulary words that we artificially split into multiple tokens. In both cases, our experiments indicate a word-level detokenization process in LLMs, which occurs mainly in the early to middle layers. Our results hint that language models hold a latent vocabulary or inner lexicon, which they access to identify words from token sequences. [^2] This process might resemble the mental lexicon in humans (Aitchison, 2012; Marslen-Wilson et al., 1994).
@@
We begin by examining if internal representations of token sequences reflect whether a sequence of tokens comprises a word (Section 3). We probe the model’s hidden representations (Conneau et al., 2018) of both multi-token real English words and gibberish nonwords (Frisch et al., 2000). We observe that the representations of words and nonwords substantially diverge in middle layers—a simple k-nearest neighbors classifier achieves a 89% accuracy on the task of discriminating between both groups. Overall, these results suggest that models hold a concept of recognized words.
@@
We next explore the mechanism through which models reconstruct cohesive word representations from sub-word tokens (Section 4). To do so, we use techniques that interpret a token’s hidden states and decode them into natural language (Belrose et al., 2023; Ghandeharioun et al., 2024). We find that for both multi-token and single-token words, in most cases, the last token can be decoded as the full word after being processed by 3–5 layers, with some words requiring up to 15 layers. Interestingly, 23% of the multi-token words are never successfully decoded from their last token’s hidden states, hinting this inner lexicon does not cover all words.
@@
Figure 1: Left: The sub-word detokenization process in LLMs. From bottom to top: (a) Tokenization and Embedding: The input string is tokenized using a sub-word tokenizer (e.g., BPE) and converted into vector embeddings; (b) Token Aggregation: the attention mechanism is used to merge sub-word tokens into the final sub-word token; (c) Word Retrieval: The model retrieves the full word representation from an internal dictionary in its feedforward (FFN) layers. This representation is added to the residual stream, until it takes over the word’s hidden representation.
@@
Right: When taking this hidden representation and patching into another prompt, the model interprets it as the original word. In this example, the model is prompted to repeat the (single vector) hidden representation of the word “unhappiness” (originally represented as three tokens), and is able to “understand” it by regenerating the original three tokens.
@@
We then turn to explore how models assemble these full word representations (Section 5). We first interpret the feedforward network (FFN) updates in the vocabulary space (Geva et al., 2021; Todd et al., 2024). We show that in 85% of the evaluated words, an FFN update promoting the full word’s concept is written to the last token’s residual stream (Geva et al., 2022b; Merullo et al., 2024). Importantly, this happens in the layers before the word’s representation emerges, indicating that models retrieve the reconstructed word representation from their FFN weights. Second, we compare attention patterns between standard single-token words to how multi-token words attend to their preceding sub-word tokens. We find that the last token in multi-token words attends significantly more to its previous (sub-word) tokens in layers 1–2 compared to single-token words, and significantly less in the following layers. This suggests that models initially aggregate information from previous tokens, and later almost ignore them.
@@
Our findings have concrete applications (Section 6); they allow us to expand the model input and output BPE vocabulary with the fused representation of multi-token words found in our experiments. Importantly, this expansion requires no update to model parameters. Applying our approach to multi-token words found in English Wikipedia data (Merity et al., 2017), we find that the model successfully uses the new vocabulary entries both as inputs and during generation; its language model performance is maintained, and even slightly improves. This demonstrates the potential to dramatically reduce both input and output sequence length, and inference costs accordingly, especially in languages where the ratio of tokens per word is high (Ahia et al., 2023; Petrov et al., 2024).
@@
Overall, our results establish that word-level detokenization is a core process in LLMs, and provide evidence of how it unfolds across model layers. Beyond improving our understanding of the internal mechanisms driving LLMs, our work lays the foundation for practical applications, particularly in terms of optimizing token management and reducing computational costs. [^3] Our code will be released here: https://github.com/schwartz-lab-NLP/Tokens2Words
@@ 2 Related Work
Sub-word tokenization algorithms (Wu, 2016; Kudo, 2018) are the standard for pre-processing text in modern LLMs. The most widely used method is byte pair encoding (BPE; Sennrich et al., 2016), which keeps frequent words intact and splits rare words into multiple sub-words. Recent studies propose ways to improve tokenization to ease word reconstruction (Provilkov et al., 2020; Hofmann et al., 2022; Yehezkel & Pinter, 2023; Bauwens & Delobelle, 2024) or analyze how tokenization affects model performance (Bostrom & Durrett, 2020; Church, 2020; Klein & Tsarfaty, 2020; Zouhar et al., 2023; Schmidt et al., 2024). To the best of our knowledge, we are the first to thoroughly investigate how LLMs internally reconstruct word representations.
@@
Recent studies suggested that early LLM layers integrate local context to map raw token embeddings into coherent concepts or entities, a process called detokenization, but their observations were derived from specific case studies (Elhage et al., 2022; Lad et al., 2024). Prior work probed transformer models to explore how predictions are constructed over the course of the model, showing early layers provide local syntactic information (Tenney et al., 2019; Vulić et al., 2020; Durrani et al., 2020). Ferrando & Voita (2024) analyzed token attributions in Llama2 (Touvron et al., 2023), and observed attention heads that promote sub-word merging. Our work focuses on word-level detokenization, and goes beyond previous efforts to provide an in-depth analysis of how word representations are assembled from multiple tokens.
@@
Recent methods for interpreting the intermediate states of LLMs draw on a residual stream perspective of transformer architectures (Vaswani et al., 2017): the hidden state acts as an information stream along the layers, from which information is read at each layer, and new information is added through residual connections (Elhage et al., 2022). Thus, hidden states in any layer can be projected into the model’s vocabulary space, treating the hidden state as if it were the output of the last layer (nostalgebraist, 2020; Dar et al., 2023; Belrose et al., 2023; Yom Din et al., 2024). Similarly, Ghandeharioun et al. (2024) proposed to decode information from hidden representations into natural language, by patching (Zhang & Nanda, 2024) it into a prompt that encourages the model to verbalize the encoded information. We use both approaches to inspect how token representations evolve across layers.
@@
Our hypothesis that models maintain an inner lexicon is motivated by recent work showing feedforward networks (FFN) layers in transformers act as key-value memories that encode factual and linguistic knowledge (Geva et al., 2021; Meng et al., 2022; Dai et al., 2022). Particularly, FFN layers were shown to enrich entity tokens with associated information (Meng et al., 2022; Geva et al., 2023) and promote relevant concepts in vocabulary space to build up predictions (Geva et al., 2022a), similar to our finding that word representations are pulled from FFN layers before emerging in the hidden state of the word’s last token. Concurrently to our work, Feucht et al. (2024) showed that models tend to “forget” the contents of early tokens in the last token representations of multi-token words, but not in single-token words. They also hypothesize the existence of an implicit vocabulary in LLMs, but do not further analyze the detokenization process. Combined with our results, their findings suggest that full word information retrieved from FFN layers “deletes” the existing representation rather than enriching it.
@@ 3 A Motivating Observation: LLMs Can Tell Words from Non-Words
One of our key hypotheses in this work is that LLMs hold an internal lexicon of words, which is different from the BPE lexicon. As a result, we begin by asking whether LLMs, when processing a sequence of tokens, capture some notion of whether or not this sequence forms a word.
To address this question, we construct a balanced dataset containing two groups: one with real English words, and another with artificially generated, meaningless nonwords (Frisch et al., 2000). Both groups are tokenized using the Llama2 tokenizer (Touvron et al., 2023). The word dataset consists of 10,000 distinct words sampled from the Gutenberg corpus (Gerlach & Font-Clos, 2018), with 53% of the words containing two tokens, 37.3% containing three tokens, and the rest four tokens. We generate the nonwords by shuffling tokens from the word dataset, ensuring that the prefix and suffix positions align with the original tokens’ positional probabilities. This process ensures that nonwords tokens maintain position distribution properties similar to word tokens (Fig. 2(a)). We next fit a $k$-nearest neighbors ($k$-NN) classifier ($k=4$, using Euclidean distance) on the hidden states of the last tokens of both words and nonwords, for each layer of the Llama2-7B model. The training set consists of 80% of the dataset, and the remaining 20% are used for evaluation.
(a) nonwords creation process
@@
Figure 2: Our word vs. nonword probing classification experiments: (2(a)) Dataset creation process. Top: real words from the Gutenberg corpus are tokenized using the Llama2 tokenizer. Bottom: nonwords are generated by shuffling tokens while maintaining common prefixes and suffixes; (2(b)) Classification results of words vs. nonwords: accuracy increases in early layers (3–7), peaks in the middle (13), and decreases in later layers (20–32).
Fig. 2(b) shows our probing results. Our results reveal a three-stage pattern in the model’s representation of word and nonword token sequences. In the model’s first few layers, representations from both groups are relatively indistinguishable and performance is close to chance level. Then, from layers 2 to 6, a clear distinction between the two groups emerges, until the representations remain almost completely distinct in middle layers, between layers 6 and 20. At this point, the probe achieves a stable, high accuracy, peaking at 89% on layer 13. Finally, performance slightly drops after layer 20. Importantly, word and nonword representations are much more similar in the representations of earlier tokens, as the probe only reaches 61% classification accuracy. We provide further details in Appendix A, as well as an analysis of wrongly classified words and nonwords.
Our results indicate that language models internally represent words and nonwords differently. This distinction is gradually developed in the model’s early layers and maintained throughout its middle layers. These findings support the hypothesis that the model performs a detokenization process and suggests where this process occurs. Building on these results, we next investigate how sub-word tokens are combined into word representations across model layers, and explore the internal mechanisms that facilitate this transformation.
@@ 4 Extracting Word Identity from LLM Hidden States
We have so far observed that LLMs can differentiate between words and nonwords, suggesting an internal detokenization process specific to word composition. We next dig into this process, by asking whether we can directly extract word identity from the hidden states of sub-word tokens. We start by considering single-token words (Section 4.1), and then move on to multi-token words (Section 4.2).
@@ 4.1 Single-Token Words
We first consider in-vocabulary words, which are mapped to single tokens by the tokenizer. Naively, such words don’t tell us much about detokenization, as they are represented using only one token. To address this, we artificially split them into multiple sub-word tokens. E.g., we take the single-token word “cats” and split it into two tokens: “ca” and ”ts”. We hypothesize that if the model performs detokenization, it will represent the last token of the word (“ts”) similarly to the original word token (“cats”). We iterate the WikiText-103 dataset (Merity et al., 2017) and randomly split each single-token word longer than three characters into 2–5 sub-words tokens. We then feed the model with the new sequence of tokens.
@@
To measure the similarity between the representation of the final token and original word, we apply the logit lens method (nostalgebraist, 2020), an interpretability method that maps a hidden representation of a given token to the word whose vector in the output unembedding matrix is most closely aligned with the hidden representation. We note that this technique is typically used to inspect model predictions of the next token in intermediate layers. As we aim to identify how LLMs represent the current word, we use the input embeddings matrix instead.
@@
We use four LLMs: Llama2-7B, Llama3-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Yi-6B (AI et al., 2024). We report for each layer the rate of retrieval—the proportion of words for which the closest vector in the BPE space is the original (single token) word.
@@
Our results for the Llama2-7B are shown in Fig. 3(a). [^4] Results for the other models show a similar trend, see Appendix C. The results show that starting layer at 8, the hidden state of the final token is mapped to the original word with high accuracy, which peaks at more than 80% in layer 15.555Further, we find that 93.2% of all split words are correctly mapped to the original word in at least one model layer (see Appendix D). Interestingly, afterwards this accuracy starts to decline, possibly because the model transitions into representing the prediction of the next token.
@@
Our results indicate that LLMs perform a detokenization process in cases where a single-token word is split into multiple tokens. This process assigns the final token with a hidden representation similar to that of the full, single-token word. But how robust is this process? To address this question, we consider a different way of splitting a word into sub-tokens—by adding typos such as flipping two adjacent characters, deleting a character, and inserting a new character—see Appendix E for more details. Much like our previous experiment, this process also splits the word into multiple tokens, as corrupted words are very rarely found in BPE lexicons. We repeat the same experiment, aiming to study whether the model maps the last token to the original (correctly-spelled) word.
@@
Our results for Llama2-7B (Fig. 3(a)) show a similar trend, though less pronounced: the model correctly matches about 40% of the corrupted words with the original, single-token, uncorrepted word. [^6] Here, 66% of the words are correctly mapped in at least one layer. We also note that the trend across layers persists, with accuracy peaking around layer 15.
@@
Figure 3:  (3(a)) Logit lens rate of retrieval of single-token words artificially split (blue) and split due to typos (orange). In both cases, we see an increasing rate of retrieval after the 4th layer, peaking in the middle (16–17 layers) and then dropping; (3(b)) Patchscopes rate of retrieval for multi-token words. In this case results peak in the 5th layer and then start to decline.

### v3

# From Tokens to Words: On Detokenization and the Inner Lexicon of LLMs

@@ Abstract
Natural language is composed of words, but modern large language models (LLMs) process sub-words as input. A natural question raised by this discrepancy is whether LLMs encode words internally, and if so how. We present evidence that LLMs engage in an intrinsic detokenization process, where sub-word sequences are combined into coherent whole-word representations at their last token. Our experiments show that this process primarily takes place within the early and middle layers of the model. We further demonstrate its robustness to arbitrary splits (e.g., “cats” to “ca” and “ts”), typos, and importantly—to out-of-vocabulary words: when feeding the last token internal representations of such words to the model as input, it can “understand” them as the complete word despite never seeing such representations as input during training. Our findings suggest that LLMs maintain a latent vocabulary beyond the tokenizer’s scope. These insights provide a practical, finetuning-free application for expanding the vocabulary of pre-trained models. By enabling the addition of new vocabulary words, we reduce input length and inference iterations, which reduces both space and model latency, with little to no loss in model accuracy. [^1] We release our code at https://github.com/schwartz-lab-NLP/Tokens2Words.
@@ 1 Introduction
Large language models (LLMs) rely heavily on tokenization methods such as byte-pair encoding (BPE; Sennrich et al., 2016). Such methods often split words into multiple tokens, potentially disrupting their morphological structure (Batsuren et al., 2024). [^2] For instance, the word “unhappiness” might be tokenized into “un,” “h,” and “appiness” (see Fig. 1) Typos and other perturbations can also lead to large variations in the tokens that represent a word (Kaushal & Mahowald, 2022). Nonetheless, LLMs exhibit a remarkable ability to recover word meaning (Cao et al., 2023). This raises important questions about how models internally compose meaningful word representations from tokens, a process referred to as detokenization (Elhage et al., 2022; Gurnee et al., 2023).
@@
In this work, we seek to understand the detokenization mechanism in LLMs. We consider two cases: words that are not part of the model’s BPE vocabulary, and are thus split into multiple sub-words; and single-token, in-vocabulary words that we artificially split into multiple tokens. In both cases, our experiments indicate a word-level detokenization process in LLMs, which occurs mainly in the early to middle layers. Our results hint that language models hold a latent vocabulary or inner lexicon, which they access to identify words from token sequences. [^3] This process might resemble the mental lexicon in humans (Aitchison, 2012; Marslen-Wilson et al., 1994).
@@
We begin by examining if internal representations of token sequences reflect whether or not a sequence of tokens comprises a word (Sec. 3). We probe the model’s hidden representations (Conneau et al., 2018) of both multi-token real English words and gibberish nonwords (Frisch et al., 2000). We observe that the representations of words and nonwords substantially diverge in middle layers—a simple k-nearest neighbors classifier achieves a 89% accuracy on the task of discriminating between the two groups. Overall, these results suggest that models hold a concept of recognized words.
@@
We next explore the mechanism through which models reconstruct cohesive word representations from sub-word tokens (Sec. 4). To do so, we use techniques that interpret a token’s hidden states and decode them into natural language (Belrose et al., 2023; Ghandeharioun et al., 2024). We find that for both multi-token and single-token words, in most cases, the last token can be decoded as the full word after being processed by 3–5 layers, with some words requiring up to 15 layers. Interestingly, 23% of the multi-token words are never successfully decoded from their last token’s hidden states, hinting this inner lexicon does not cover all words.
@@
Figure 1: Left: The sub-word detokenization process in LLMs. From bottom to top: (a) Tokenization and Embedding: The input string is tokenized using a sub-word tokenizer (e.g., BPE) and converted into vector embeddings; (b) Token Aggregation: The attention mechanism relays information from the word’s previous tokens (“un”, “h”) into its final sub-word representation (“appiness”); (c) Word Retrieval: The model retrieves the full word representation from an implicit internal lexicon in its feedforward (FFN) layers. This representation is added to the residual stream, until it takes over the word’s hidden representation.
@@
Right: When taking this hidden representation and patching it into another prompt, the model interprets it as the original word. In this example, the model is prompted to repeat the (single vector) hidden representation of the word “unhappiness” (originally represented as three tokens), and is able to “understand” it by regenerating the original three tokens.
@@
We then turn to explore how models assemble these full word representations (Sec. 5). We first interpret the feedforward network (FFN) updates in the vocabulary space (Geva et al., 2021; Todd et al., 2024). We show that in 85% of the evaluated words, an FFN update promoting the full word’s concept is written to the last token’s residual stream (Geva et al., 2022; Merullo et al., 2024). Importantly, this happens in the layers before the full word’s representation emerges, indicating that models retrieve the reconstructed word representation from their FFN weights. Second, we compare attention patterns between standard single-token words to how multi-token words attend to their preceding sub-word tokens. We find that the last token in multi-token words attends significantly more to its previous tokens (which are sub-words of the same word) in layers 1–2 compared to single-token words (where these previous tokens represent other words), and significantly less in the following layers. This suggests that models initially aggregate information from previous tokens, and later almost ignore them. See Fig. 1 for an overview.
@@
Our findings have concrete applications (Sec. 6); they allow us to expand the model input and output BPE vocabulary with the fused representation of multi-token words found in our experiments. Importantly, this expansion requires no update to model parameters. Applying our approach to multi-token words found in English Wikipedia data (Merity et al., 2017), we find that the model successfully uses the new vocabulary entries both as inputs and during generation: its language model performance is maintained, and even slightly improves. This demonstrates the potential to dramatically reduce both input and output sequence length, and inference costs accordingly, especially in languages where the ratio of tokens per word is high (Ahia et al., 2023; Petrov et al., 2023).
@@
Overall, our results establish that word-level detokenization is a core process in LLMs, and provide evidence of how it unfolds across model layers. Beyond improving our understanding of the internal mechanisms driving LLMs, our work lays the foundation for practical applications, particularly in optimizing token management and reducing computational costs.
@@ 2 Related Work
Sub-word tokenization algorithms (Wu, 2016; Kudo, 2018) are the standard for pre-processing text in modern LLMs. The most widely used method is byte-pair encoding (BPE; Sennrich et al., 2016), which keeps frequent words intact and splits rare ones into multiple sub-words. Recent studies proposed ways to improve tokenization to consider word structure (Provilkov et al., 2020; Hofmann et al., 2022; Yehezkel & Pinter, 2023; Bauwens & Delobelle, 2024) or analyze how tokenization affects model performance (Bostrom & Durrett, 2020; Church, 2020; Klein & Tsarfaty, 2020; Zouhar et al., 2023; Schmidt et al., 2024). To the best of our knowledge, we are the first to thoroughly investigate how LLMs internally reconstruct word representations.
@@
Early LLM layers have been shown to integrate local context and map raw token embeddings into representations of concepts or entities—a process called detokenization. However, such observations were based on specific case studies (Elhage et al., 2022; Lad et al., 2024). More generally, recent work showed early layers provide local syntactic information (Tenney et al., 2019; Vulić et al., 2020; Durrani et al., 2020; Sajjad et al., 2022) or focus on extracting information from previous tokens (Ben Artzy & Schwartz, 2024). Ferrando & Voita (2024) analyzed token attributions in LLMs, observing attention heads that promote sub-word merging (Correia et al., 2019). Our work focuses on word-level detokenization, and goes beyond previous efforts to provide an in-depth analysis of how word representations are assembled from multiple tokens.
@@
Recent methods for interpreting the intermediate states of LLMs draw on a residual stream perspective: the hidden state acts as an information stream along the layers, from which information is read at each layer, and new information is added through residual connections (Elhage et al., 2022). Thus, hidden states at any layer can be projected into the model’s vocabulary space, treating the hidden state as if it were the output of the last layer (nostalgebraist, 2020; Dar et al., 2023; Belrose et al., 2023; Yom Din et al., 2024). Similarly, Ghandeharioun et al. (2024) proposed to decode information from hidden representations into natural language, by patching (Zhang & Nanda, 2024) it into a prompt that encourages the model to verbalize the encoded information. We use both approaches to inspect how token representations evolve across layers.
@@
The idea of an inner lexicon aligns with recent work showing feedforward networks (FFN) layers in transformers act as key-value memories that encode factual and linguistic knowledge (Geva et al., 2021; Meng et al., 2022b; Dai et al., 2022). Particularly, FFNs were shown to enrich entity tokens with associated information (Meng et al., 2022b; Geva et al., 2023) and promote relevant concepts in vocabulary space to build up predictions (Geva et al., 2022). Our work expands on these findings, showing that word representations are pulled from FFN layers before emerging in the hidden state of the word’s last token.
Inner lexicon structure
Inspired by studies on how concepts are encoded in LLMs (Park et al., 2024; 2025) and gradually promoted throughout their layers (Geva et al., 2022; Merullo et al., 2024), we consider the inner lexicon a “soft” lexicon, which (a) combines multiple vectors to form word representations (rather than a key-value dictionary); and (b) is not unique, i.e., a word might be stored and retrieved in more than one layer. Concurrently, Feucht et al. (2024) present evidence for an implicit vocabulary in LLMs, showing that models “forget” preceding tokens in multi-token words or multi-word entities, but remember previous tokens when processing single-token words.
@@ 3 A Motivating Observation: LLMs Can Tell Words From Non-Words
One of our key hypotheses in this work is that LLMs hold an internal lexicon of words, which is different from the BPE lexicon. We begin by asking whether LLMs, when processing a sequence of tokens, capture some notion of whether or not this sequence forms a word.
To address this question, we construct a balanced dataset containing two groups: one with real English words, and another with artificially generated, meaningless nonwords (Frisch et al., 2000). Both groups are tokenized using the Llama2 tokenizer (Touvron et al., 2023). The word dataset consists of 10,000 distinct words sampled from the Gutenberg corpus (Gerlach & Font-Clos, 2018), with 53% of the words containing two tokens, 37.3% containing three tokens, and the rest four tokens. We generate the nonwords by shuffling tokens from the word dataset, ensuring that the prefix and suffix positions align with the original tokens’ positional probabilities. For example, the token “ing”, extracted from the final position of real words in the dataset (e.g., “directing” tokenized as “direct”, “ing”), is retained as a suffix in the nonwords dataset. This process ensures that nonword tokens maintain position distribution properties similar to word tokens, preserving natural positional patterns and mitigating potential distributional biases (Fig. 2(a)). We next apply a $k$-nearest neighbors ($k$-NN) probing classifier ($k=4$, using Euclidean distance) on the hidden states of the last tokens of both words and nonwords, for each layer of the Llama2-7B model. The training set consists of 80% of the dataset, and the remaining 20% are used for evaluation.
(a) Nonword creation process
@@
Figure 2: Our word vs. nonword probing classification experiments. (2(a)) Dataset creation process. Top: words from the Gutenberg corpus are tokenized using the Llama2 tokenizer. Bottom: nonwords are generated by shuffling tokens while maintaining their positions within the word; (2(b)Classification results of words vs. nonwords. Using the last token shows strong results (reaching up to 90% accuracy), which sharply rise in early layers (3–7), peak in the middle (13), and decrease in later layers (20–32). Using the penultimate token shows substantially lower scores, suggesting that the high classification accuracy is specifically tied to the presence of a complete word rather than token co-occurrence patterns.
Our results (Fig. 2(b), blue) reveal a three-stage pattern in the model’s representation of word and nonword token sequences. In the model’s first few layers, representations from both groups are relatively indistinguishable and accuracy is close to chance level. Then, from layers 2 to 6, a clear distinction between the two groups emerges, until the representations are almost completely separate in middle layers, between layers 6 and 20. At this point, the probe achieves a stable, high accuracy, peaking at 89% on layer 13. Finally, accuracy slightly drops after layer 20.444We perform a similar experiment using a dataset of morphologically plausible nonwords (ARC Nonword Database; Rastle et al. 2002), which shows a similar trend. See App. A.
Our results indicate the LLMs can distinguish between words and nonwords. But this distinction can be attributed to the distributional properties of words: it might be the case that models are recognizing common sequences of tokens, rather than identifying whole words. To test this hypothesis, we repeat the same experiment, this time using the penultimate token representation, for words three tokens or longer. By definition, such words also as frequently co-occur with the initial sub-word tokens in their words as the final tokens. Our results (Fig. 2(b), orange line) show that the probe only reaches 61% classification accuracy, indicating that the high classification accuracy does not stem from token co-occurrence, but is tied to the presence of a whole word. See App. A for an analysis of misclassified words and nonwords.
Overall, our results show that language models internally represent words and nonwords differently. This distinction is gradually developed in the model’s early layers and maintained throughout its middle layers. These findings support the hypothesis that the model performs a detokenization process and suggests where this process occurs. Building on these results, we next investigate how sub-word tokens are combined into word representations across model layers, and explore the internal mechanisms that facilitate this transformation.
@@ 4 Extracting Word Identity From LLM Hidden States
We have so far observed that LLMs can differentiate between words and nonwords, suggesting an internal detokenization process specific to word composition. We next dig into this process, by asking whether we can directly extract word identity from the hidden states of sub-word tokens. We start by considering single-token words (Sec. 4.1), and then move on to multi-token words (Sec. 4.2).
@@ 4.1 Single-Token Words
We first consider in-vocabulary words, which are mapped to single tokens by the tokenizer. Naively, such words don’t tell us much about detokenization, as they are represented using only one token. To address this, we artificially split them into multiple sub-word tokens. For example, we take the single-token word “cats” and split it into two tokens: “ca” and ”ts”. We hypothesize that if the model performs detokenization, it will represent the last token of the word (“ts”) similarly to the original word token (“cats”). We iterate the WikiText-103 dataset (Merity et al., 2017) and randomly split each single-token word longer than three characters into 2–5 sub-words tokens. We then feed the model the new sequence of tokens preceded by the last 100 tokens that came before each split word in the original text as context. [^5] We observe similar trends, in this and in other experiments in this section, when passing the split words without their context, albeit with slightly lower rates of retrieval.
@@
To measure the similarity between the representation of the final token and original word, we apply the logit lens method (nostalgebraist, 2020), an interpretability method that maps a hidden representation of a given token to the word whose vector in the output unembedding matrix is most closely aligned with the hidden representation. We note that this technique is typically used to inspect model predictions of the next token in intermediate layers. As we aim to identify how LLMs represent the current word, we use the input embeddings matrix rather than the unembedding one. [^6] We also compute a second measure, cosine similarity, to compare between the hidden representations and input embeddings. We find similar results to those obtained using logit lens. For details, see App. F.
@@
We study four LLMs: Llama2-7B, Llama3-8B (Dubey et al., 2024), Mistral-7B (Jiang et al., 2023), and Yi-6B (AI et al., 2024). For each layer, we report the rate of retrieval—the proportion of words for which the closest vector in the vocabulary space is the original (single token) word.
@@
Our results for the Llama2-7B are shown in Fig. 3(a). [^7] Results for the other models (App. C) show a similar trend. Starting at layer 8, the hidden state of the final token is mapped to the original word with high accuracy, which peaks at more than 80% in layer 15. Interestingly, this accuracy then starts to decline, possibly because the model transitions into representing the prediction of the next token. We also find that 93.2% of all split words are correctly mapped to the original word in at least one model layer (see Sec. C.1).
@@
Our results indicate that LLMs perform a detokenization process in cases where a single-token word is split into multiple tokens. This process assigns the final token with a hidden representation similar to that of the full, single-token word. But how robust is this process? To address this question, we consider a different way of splitting a word into sub-tokens—adding typos. For each word, we randomly flip two adjacent characters, delete a character, or insert a new character (see App. D for more details). Much like our previous experiment, this process splits the word into multiple tokens, as corrupted words are rarely found in BPE lexicons. We repeat the same experiment as above, aiming to study whether models map the last token to the original (correctly-spelled) word.
@@
Our results for Llama2-7B (Fig. 3(a)) show a similar trend, though less pronounced: the model correctly matches about 40% of corrupted words with the original, single-token, uncorrupted word. [^8] Here, 66% of the words are correctly mapped in at least one layer. We also note that the trend across layers persists, with accuracy peaking around layer 15.
@@
Figure 3:  (3(a)) Logit lens rate of retrieval of single-token words artificially split (blue line) and split due to typos (orange). In both cases, we see an increasing rate of retrieval after the 4th layer, peaking in the middle (16–17 layers) and then dropping; (3(b)) Patchscopes rate of retrieval for multi-token words. In this case results peak in the 5th layer and then start to decline.